# Kac–Moody Symmetric Spaces: arbitrary symmetrizable complex or almost split real type

### v4

@@ 5.1 Global Action
From 4.12 it follows that $G$ acts transitively on standard flats. Since all standard flats are maximal flats and all maximal flats according to 4.10 are Euclidean in our context, all assumptions of [FHHK20, Proposition 2.33] are fulfilled. ∎
@@
To do this, we need to recall some things from Section 3.2. First, the map $\eta$ from the Tits functor, further the fact, that an element of the root group $U_{\alpha}$ is given by $x_{\alpha}(r)$ for $r\in\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and $\alpha\in\Delta^{re}$. And lastly, that the torus of a Kac–Moody group $T=\mathfrak{T}_{\Lambda}(\mathbb{K})$ is isomorphic to $\operatorname{Hom}_{grp}(\Lambda,\mathbb{K}^{\times})$. Hence we can define the map
@@
    $\left(\nicefrac{{G^{\dagger}}}{{K^{\dagger}}},\mu\mid_{G^{\dagger}}\right)$
@@ 6 Local Action
This section is primarily based on [FHHK20, Section 6.14]. We will sometimes refer to statements from [FHHK20] without justifying why, since the statements we refer to do not take into account the underlying field or type of the Kac–Moody group. If there is a statement that does take this into account and needs further explanation before we apply it in our context, we will point it out.
@@ 6.1 Comparing the Weyl Groups
$\nicefrac{{N_{G^{\dagger}}\left(T^{\dagger}\right)}}{{T^{\dagger}}}\cong\mathcal{W}\cong\nicefrac{{N_{G}(T)}}{{T}}.$ (9)
@@
and recall from 3.14 that $M\subset K$. In the proof of [FHHK20, Lemma 3.27 (ii)], using the Iwasawa decomposition and the refined Birkhoff decomposition, it is shown that $N_{G}(T)=N_{K}(T)T$ holds. Note that the refined Birkhoff decomposition can be used since the $BN$-pair obtained from $G$ is refined according to [Rém02, Theorem 1.5.4]. The restriction $\pi\mid_{N_{K}(T)}$ is therefore still surjective. From [FHHK20, Lemma 3.27 (ii)] it now follows that
@@
and therefore the kernel of $\pi\mid_{N_{K}(T)}$ is equal to $M$. Thus
@@
$W\left(G\curvearrowright G/K\right)\cong\mathcal{W}.$
@@ 6.2 A Local to Global Approach
[…]
@@
The Chevalley involution $\Theta$ acts on the standard flat by inversion and so it corresponds to $\mathbb{Z}/2\mathbb{Z}$. Note that the semidirect product with $\Theta$ vanishes in the decomposition of the stabilizer in $\operatorname{Aut}_{\text{eff}}(G^{\dagger})$ as one considers the action on $A^{\dagger}K^{\dagger}$ and by the proof of [FHHK20, Proposition 6.5] it is known that the Chevalley involution commutes with $\operatorname{Aut}(\Gamma)$ and it can be concluded that $\Theta$ also commutes with the inner automorphisms. In fact, since one considers the stabilizers of the standard flat $A^{\dagger}K^{\dagger}$, one knows that $\Theta$ commutes with the diagram automorphism and normalizes the left multiplication. If we denote the left multiplication with $g\in\operatorname{Stab}_{G^{\dagger}}(e,A^{\dagger}K^{\dagger})$ by $l(g)$, then the following is true
@@
$\operatorname{Stab}_{G^{\dagger}}(e,A^{\dagger}K^{\dagger})\cong N_{K^{\dagger}}(T^{\dagger})\quad\text{and}\quad\operatorname{Fix}_{G^{\dagger}}(e,A^{\dagger}K^{\dagger})\cong M^{\dagger}.$
@@
/StabG(e,AK)/CK(G)FixG(e,AK)/CK(G) $\cong\nicefrac{{\operatorname{Stab}_{G^{\dagger}}(e,A^{\dagger}K^{\dagger})}}{{\operatorname{Fix}_{G^{\dagger}}(e,A^{\dagger}K^{\dagger})}}$
@@ 7 Comparing Symmetric Spaces
Regarding the case of generalizing the topological Iwasawa decomposition to complex split Kac–Moody groups, there is now a way to do this thanks to [HK23, Corollary B.8]. The quoted statement allows us to use the same methods as in [FHHK20] to prove the topological Iwasawa decomposition in the complex case, see 3.20.
@@ 8.1 Real Forms of Kac–Moody Groups
[…]
  • An $\mathbb{R}$-functor $F$ is called a $\mathbb{R}$-form of a $\mathbb{C}$-functor $H$, if there is a functorial isomorphism $F\mid_{\mathbb{C}}\cong H$, i.e. the functor $F$ applied to the field $\mathbb{C}$ is functorial isomorphic to the functor $H$.
@@
[…]
@@ 8.2 The Galois Descent
To make the following terms well-defined, we need to define a triple $(A,C,-C)$, where $A$ is an apartment and $C,-C$ are opposite chambers in $A$. This is similar to defining a standard torus and defines the geometric realization. In our context, it is only necessary that we obtain the following when applying the theory of [Rém02, Chapter 12]:
@@
    Denote by $\musNaturalre\Delta^{\musNatural}_{\mathrm{re}}$ the set of all real $\mathbb{R}$-roots with respect to $(A,C,-C)$ (cf. [Rém02, 5.5.1]). Denote by $\musNatural\Pi^{\musNatural}$ the set of simple roots of $\musNaturalre\Delta^{\musNatural}_{\mathrm{re}}$.
@@ 8.3 Relative Kac–Peterson Topology
$\tau_{\text{MB}}^{\operatorname{as}}=\tau_{\text{MB}}\mid_{G(\mathbb{R})},$ (11)
where $\tau_{\text{MB}}\mid_{G(\mathbb{R})}$ means the subspace topology induced by $\tau_{\text{MB}}$.
@@
    $\tau_{\text{MB}}^{\operatorname{as}}=\tau_{\text{MB}}\mid_{G(\mathbb{R})}=\tau_{\text{Prod}}\mid_{G(\mathbb{R})}=\tau_{\text{KP}}\mid_{G(\mathbb{R})}$,
@@ 9 Symmetric Spaces of Almost Split Kac–Moody Groups
$\left(\Theta\circ\widehat{\sigma}\right)(x_{\alpha}(t))$ $=\Theta(x_{\overline{\sigma}(\alpha)}(\overline{t}))$
@@
$\left(\widehat{\sigma}\circ\Theta\right)(x_{\alpha}(t))$ $=\widehat{\sigma}\left(x_{-\alpha}(\overline{t})\right)$
@@
[…]
@@
$\left(\Theta\circ\overline{\sigma}\right)\left(r^{v_{i}}\right)$ $=\Theta\left(\overline{r}^{v_{\overline{\sigma}(i)}}\right)=r^{-v_{\overline{\sigma}(i)}},$
@@
$\left(\overline{\sigma}\circ\Theta\right)\left(r^{v_{i}}\right)$ $=\overline{\sigma}\left(\overline{r}^{-v_{i}}\right)=r^{-v_{\overline{\sigma}(i)}}.$
@@
$\Theta\left(G^{\Gamma}\right)=\Theta(G)^{\Gamma},$
@@
[…]
@@ 9.1 Flats in Almost Split Kac–Moody Symmetric Spaces
The next step is to analyze the image of $\musNatural\exp_{\operatorname{KM}}(\mathfrak{a}^{\musNatural})$. Therefore, let
@@
$\overline{\sigma}(\exp_{\operatorname{KM}}(h))$ $=\overline{\sigma}\left(\prod_{i=1}^{2n-l}\left(e^{r_{i}}\right)^{v_{i}}\right)=\prod_{i=1}^{2n-l}\sigma(e^{r_{i}})^{v_{\overline{\sigma}(i)}}=\prod_{i=1}^{2n-l}(e^{\overline{r_{i}}})^{v_{\overline{\sigma}(i)}}$
$\exp_{\operatorname{KM}}(\overline{\sigma}(h))$ $=\exp_{\operatorname{KM}}\left(\sum_{i=1}^{2n-l}\overline{r_{i}}v_{\overline{\sigma}(i)}\right)=\prod_{i=1}^{2n-l}(e^{\overline{r_{i}}})^{v_{\overline{\sigma}(i)}}.$
@@
Note that $\mathfrak{a}^{\musNatural}\cong\mathbb{R}^{m}$, where $m<2n-l$, but the actual size of $m$ is not clear. It depends on how exactly the torus of $G(\mathbb{R})$ looks and splits.
@@
Recall from 3.20 that an element $g\in AK$ can be written as $ak$ for a unique $a\in A$ and a unique $k\in K$. Assume that $ak\in AK\cap G(\mathbb{R})\subseteq G$ and let $\overline{\sigma}$ a rectified Galois automorphism. Then we have $\overline{\sigma}(A)\subseteq A$ and $\overline{\sigma}(K)\subseteq K$. From the condition (8.2) it further follows that

### v5

@@ 5.1 Global Action
From 4.12 it follows that $G$ acts transitively on standard flats. Since all standard flats are maximal flats and all maximal flats according to 4.10 are Euclidean in our context, all assumptions of [FHHK20, Proposition 2.33] are satisfied. ∎
@@
To do this, we need to recall some facts from Section 3.2. Firstly, the map $\eta$ from the Tits functor; secondly, the fact that an element of the root group $U_{\alpha}$ is given by $x_{\alpha}(r)$ for $r\in\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and $\alpha\in\Delta^{re}$; and, lastly, that the torus of a Kac–Moody group $T=\mathfrak{T}_{\Lambda}(\mathbb{K})$ is isomorphic to $\operatorname{Hom}_{grp}(\Lambda,\mathbb{K}^{\times})$. Hence we can define the map
@@
    $\left(\nicefrac{{G^{\dagger}}}{{K^{\dagger}}},\mu\hskip-3.0pt\mid_{G^{\dagger}}\right)$
@@ 6 Local Action
This section is primarily based on [FHHK20, Section 6.14]. We will sometimes refer to statements from [FHHK20] without any justification when the statements we are referring to do not depend on the underlying field or the type of the Kac–Moody group. We will point out statements that require further explanation before being applicable into our context.
@@ 6.1 Comparing the Weyl Groups
$\nicefrac{{N_{G^{\dagger}}\hskip-1.5pt\left(T^{\dagger}\right)}}{{T^{\dagger}}}\cong\mathcal{W}\cong\nicefrac{{N_{G}(T)}}{{T}}.$ (9)
@@
and recall from 3.14 that $M\subset K$. In the proof of [FHHK20, Lemma 3.27 (ii)], using the Iwasawa decomposition and the refined Birkhoff decomposition, it is shown that $N_{G}(T)=N_{K}(T)T$ holds. Note that the refined Birkhoff decomposition can be used since the $BN$-pair obtained from $G$ is refined according to [Rém02, Theorem 1.5.4]. The restriction $\pi\hskip-3.0pt\mid_{N_{K}(T)}$ is therefore still surjective. From [FHHK20, Lemma 3.27 (ii)] it now follows that
@@
and therefore the kernel of $\pi\hskip-3.0pt\mid_{N_{K}(T)}$ is equal to $M$. Thus
@@
$W\hskip-2.0pt\left(G\curvearrowright G/K\right)\cong\mathcal{W}.$
@@ 6.2 A Local to Global Approach
$W(\operatorname{Aut}(X^{\dagger})\curvearrowright X^{\dagger})$$\operatorname{GL}\left(p,F,{F^{\text{sing}}}(p)\right)$$\left(\mathcal{W}\rtimes\operatorname{Aut}(\Gamma)\right)\times\mathbb{Z}/2\mathbb{Z}$$\mathbb{R}_{>0}\times\left(\mathcal{W}\rtimes\operatorname{Aut}(\mathcal{W},S)\right)\times\mathbb{Z}/2\mathbb{Z}$$\kappa$$\iota\cong\cong$
@@
The Chevalley involution $\Theta$ acts on the standard flat by inversion and so it corresponds to $\mathbb{Z}/2\mathbb{Z}$. Note that the semidirect product with $\Theta$ vanishes in the decomposition of the stabilizer in $\operatorname{Aut}_{\text{eff}}(G^{\dagger})$ as one considers the action on $A^{\dagger}K^{\dagger}$ and by the proof of [FHHK20, Proposition 6.5] it is known that the Chevalley involution commutes with $\operatorname{Aut}(\Gamma)$ and it can be concluded that $\Theta$ also commutes with the inner automorphisms. In fact, since one considers the stabilizers of the standard flat $A^{\dagger}K^{\dagger}$, one knows that $\Theta$ commutes with the diagram automorphism and normalizes the left multiplication. If we denote the left multiplication with $g\in\operatorname{Stab}_{G^{\dagger}}\hskip-2.0pt(e,A^{\dagger}K^{\dagger})$ by $l(g)$, then the following is true
@@
$\operatorname{Stab}_{G^{\dagger}}\hskip-2.0pt(e,A^{\dagger}K^{\dagger})\cong N_{K^{\dagger}}(T^{\dagger})\quad\text{and}\quad\operatorname{Fix}_{G^{\dagger}}\hskip-2.0pt(e,A^{\dagger}K^{\dagger})\cong M^{\dagger}.$
@@
/StabG(e,AK)/CK(G)FixG(e,AK)/CK(G) $\cong\nicefrac{{\operatorname{Stab}_{G^{\dagger}}\hskip-2.0pt(e,A^{\dagger}K^{\dagger})}}{{\operatorname{Fix}_{G^{\dagger}}\hskip-2.0pt(e,A^{\dagger}K^{\dagger})}}$
@@ 7 Comparing Symmetric Spaces
Regarding the case of generalizing the topological Iwasawa decomposition to complex split Kac–Moody groups, there is now a way to do this thanks to [HK23, Corollary B.8]. The quoted statement allows us to use the same methods as in [FHHK20] to prove the topological Iwasawa decomposition in the complex case, see 3.19.
@@ 8.1 Real Forms of Kac–Moody Groups
  • Let $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$. Call a functor $F\colon\mathbb{K}\text{-alg}\to\textbf{Grp}$ a $\mathbb{K}$-functor.
  • An $\mathbb{R}$-functor $F$ is called a $\mathbb{R}$-form of a $\mathbb{C}$-functor $H$, if there is a functorial isomorphism $F\hskip-3.0pt\mid_{\mathbb{C}}\cong H$, i.e. the functor $F$ applied to the field $\mathbb{C}$ is functorial isomorphic to the functor $H$.
@@
  • The adjoint representation $\operatorname{Ad}\colon\mathcal{G}\to\operatorname{Aut}(\mathcal{U})$, where $\operatorname{Aut}(\mathcal{U})\colon\mathbb{Z}\text{-alg}\to\textbf{Grp}$ is a functor, is equivariant with respect to the Galois action.
@@ 8.2 The Galois Descent
To make the following terms well-defined, we need to define a triple (A,C,C)(A,C,-C), where $A$ is an apartment and C,$C$,-C are opposite chambers in $A$. This is similar to defining a standard torus and defines the geometric realization. In our context, it is only necessary that we obtain the following when applying the theory of [Rém02, Chapter 12]:
@@
    Denote by $\musNatural\Delta^{\musNatural}_{\mathrm{re}}$ the set of all real $\mathbb{R}$-roots with respect to (A,C,C)(A,C,-C) (cf. [Rém02, 5.5.1]). Denote by $\musNatural\Pi^{\musNatural}$ the set of simple roots of $\musNatural\Delta^{\musNatural}_{\mathrm{re}}$.
@@ 8.3 Relative Kac–Peterson Topology
$\tau_{\text{MB}}^{\operatorname{as}}=\tau_{\text{MB}}\hskip-3.0pt\mid_{G(\mathbb{R})},$ (11)
where $\tau_{\text{MB}}\hskip-3.0pt\mid_{G(\mathbb{R})}$ means the subspace topology induced by $\tau_{\text{MB}}$.
@@
    $\tau_{\text{MB}}^{\operatorname{as}}=\tau_{\text{MB}}\hskip-3.0pt\mid_{G(\mathbb{R})}=\tau_{\text{Prod}}\hskip-3.0pt\mid_{G(\mathbb{R})}=\tau_{\text{KP}}\hskip-3.0pt\mid_{G(\mathbb{R})}$,
@@ 9 Symmetric Spaces of Almost Split Kac–Moody Groups
$\left(\Theta\circ\widehat{\sigma}\right)\hskip-2.0pt(x_{\alpha}(t))$ $=\Theta(x_{\overline{\sigma}(\alpha)}(\overline{t}))$
@@
$\left(\widehat{\sigma}\circ\Theta\right)\hskip-2.0pt(x_{\alpha}(t))$ $=\widehat{\sigma}\left(x_{-\alpha}(\overline{t})\right)$
@@
It remains to show, that $\Theta$ and $\widehat{\sigma}$ commutes on $T$ and since every diagonal automorphism is the identity on $T$, we conclude that $\widehat{\sigma}$ acts as $\overline{\sigma}$ on $T$. The action of $\overline{\sigma}$ on $T$ is given by Section 8.1 3. Now given a generating set $\{r^{v_{i}}\mid r\in\mathbb{C}^{\times}\,,\;1\leq i\leq 2n-l\}$ of $T$. Then we calculate for any $\alpha\in\Delta^{\mathrm{re}}$
@@
$\left(\Theta\circ\overline{\sigma}\right)\hskip-2.0pt\left(r^{v_{i}}\right)$ $=\Theta\left(\overline{r}^{v_{\overline{\sigma}(i)}}\right)=r^{-v_{\overline{\sigma}(i)}},$
@@
$\left(\overline{\sigma}\circ\Theta\right)\hskip-2.0pt\left(r^{v_{i}}\right)$ $=\overline{\sigma}\left(\overline{r}^{-v_{i}}\right)=r^{-v_{\overline{\sigma}(i)}}.$
@@
$\Theta\hskip-2.0pt\left(G^{\Gamma}\right)=\Theta(G)^{\Gamma},$
@@
Note, that $\tau^{\operatorname{as}}$ equals $\tau$ since $G(\mathbb{R})$ is a subgroup of $G$. Moreover, since $G(\mathbb{R})$ is closed with respect to $\Theta$ and $\tau^{\operatorname{as}}$, the map $\mu^{\operatorname{as}}$ satisfies all properties of an abstract symmetric space (cf. 2.1). Even the fourth property $\textbf{S.4}_{\text{global}}$ is true, because
@@ 9.1 Flats in Almost Split Kac–Moody Symmetric Spaces
Note, that $\musNatural\mathfrak{a}^{\musNatural}$ is exactly the fixed point set of the real form $\mathfrak{a}$ under the Galois action. Recall, that $G(\mathbb{R})\leq G$ and by $\musNatural𝔞\mathfrak$$^{\musNatural}\subseteq\mathfrak{a}$ we can use the exponential map given in Section 3.2: restricting the map onto $\musNatural\mathfrak{a}^{\musNatural}$ leads to
@@
$\overline{\sigma}(\exp_{\operatorname{KM}}(h))$ $=\overline{\sigma}\hskip-2.0pt\left(\prod_{i=1}^{2n-l}\left(e^{r_{i}}\right)^{v_{i}}\right)=\prod_{i=1}^{2n-l}\sigma(e^{r_{i}})^{v_{\overline{\sigma}(i)}}=\prod_{i=1}^{2n-l}(e^{\overline{r_{i}}})^{v_{\overline{\sigma}(i)}}$
$\exp_{\operatorname{KM}}(\overline{\sigma}(h))$ $=\exp_{\operatorname{KM}}\hskip-2.0pt\left(\sum_{i=1}^{2n-l}\overline{r_{i}}v_{\overline{\sigma}(i)}\right)=\prod_{i=1}^{2n-l}(e^{\overline{r_{i}}})^{v_{\overline{\sigma}(i)}}.$
@@
[…]
@@
Recall from 3.19 that an element $g\in AK$ can be written as $ak$ for a unique $a\in A$ and a unique $k\in K$. Assume that $ak\in AK\cap G(\mathbb{R})\subseteq G$ and let $\overline{\sigma}$ a rectified Galois automorphism. Then we have $\overline{\sigma}(A)\subseteq A$ and $\overline{\sigma}(K)\subseteq K$. From the condition (8.2) it further follows that